I have a problem with applying the order of operations here are some problems that i have tried: 4^3-5(7-[1])
7(3-9)+12/(-4)

Answer by CubeyThePenguin(3113)

4^3 - 5(7 - 1)
= 64 - 5(6)
= 64 - 30 = 34

7(3 - 9) + 12/(-4)
= 7(-6) + -3
= -42 - 3 = -45